The function f(x) = 5-5x^4 has an absolute value of what and this occurs at x equals
1
Expert's answer
2020-11-24T14:00:30-0500

x^4 has minimum value at x=0, -x^4 in point 0 has absolut point, means "f(x) = 5-5x^{4}" has absolute value at x=0 f(0)=5